question 1:what are all the measurements of a cone we can solve for in math?, pls list all

question 2: after you list all measurements we can solve for cone shape, pls list what given values do we need to each of the measurements to be able to answer it
Answer #1

Cone V A Cone is a solid that has a circular base and a single vertex. If the vertex is over the center of the base, it is caCone length of arc =circumference v of base Formation of a cone : · A Cone is formed by bringing together the two radii of thCone Curved(lateral) Surface Area. The diameter of a cone is 14 cms. Its slant height is locms. Find L.S.A l v=L(P) = 7cm CurCone Total Surface Area Tirl Total surface area = L. S. A+ Area of base = url + 11r2 Ir( T+7) = Tr( v r2 + H2+r) The height oCone Volume Volume of cone = 1/3( volume of cylinder) = 1/3 (II r2h) cubic units. The area of the base of cone is 78.5cm. If